OpenStudy (anonymous):

Original problem statement: Each valve A, B, and C, when open, releases water into a tank at its own constant rate. With all three valves open, the tank fills in 1 hour, with only valves A and C open it takes 2 hours, and with only valves B and C open it takes 3 hours. How long will it take to fill the tank with only valves A and B open? Analysis: Three valves with flow rates (vol / time) of A, B, C. Thus, in a time t, the A valve delivers a volume of At. Let the total volume of the tank equal 1 in our units. In each case below, we start at t = 0 with the tank empty.

OpenStudy (anonymous):

With all 3 valves open, the tank is filled in time t = 1, so we have: A + B + C = 1 With just A and B open, the tank is filled at time t = 2, so we have: 2A + 2B = 1 And with valves B and C open, the tank is filled at time t = 3. 3B + 3C = 1 My solution is: A = 2/3 B = -1/6 C = 1/2

OpenStudy (anonymous):

@lgbasallote original problem is here: http://openstudy.com/users/azntiger627#/updates/50171993e4b04dfc808b5196 Can you please explain your answer more, in particular use of the reciprocal.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Having trouble with links. Your analysis is above. Original problem is http://openstudy.com/users/azntiger627#/updates/50170e92e4b04dfc808b485d

OpenStudy (anonymous):

Guess I forgot to do the last part: At + Bt = 1 2/3 - 1/6 t = 1/2 t = 1 t = 2

OpenStudy (anonymous):

I see a problem with this.